Find the present value of a simple perpetuity of P10,000 payable semi-annually if the interest rate is 2% per six months and the first payment is due (a) six months from now and (b) now.


1
Expert's answer
2021-06-01T12:33:13-0400

a.

Semi-annual perpetuity=P10,000

Semi-annual interest rat=0.02

So, present value of perpetuity"=\frac{semi-annual \space perpetuity}{interest }"


"rate=\frac{P10000}{0.02}=P500000"


b.

  If first payment is to be received now,

present value of perpetuity=installment amount+ present value of ordinary perpetuity

"=P10,000+P500000= P510000"